OpenStudy (anonymous):

***Help Please?!?*** ***Medal AND Fan Reward*** Two forces of 5 lb. and 14 lb. act on a body at right angles to each other. Find their resultant. A)13 B)14 C)15

OpenStudy (anonymous):

\[|F_R|=\sqrt{5^2+14^2}=\sqrt{221}=?\]

OpenStudy (anonymous):

pythagoras theorem..about 15

OpenStudy (anonymous):